# Revisiting GW150914 with a non-planar, eccentric waveform model

### v1

@@ 1 Introduction
where $|\bm{\chi}_{i}^{\perp}|$ are the projections of the spin vectors onto the plane perpendicular to $\bm{L}$. The waveform polarizations $h_{+}$, $h_{\times}$ are decomposed in terms of the $s=-2$ spin-weighted spherical harmonics ${}_{-2}Y_{lm}$ as
@@ 2 Modeling GWs from BBHs on generic orbits
Figure 2: Comparison of the $(2,2)$ and $(2,1)$ modes of the TEOBResumS-Dalí model against the SXS:BBH:1389 simulation, which is a spin-precessing simulation with mass ratio $q=1.63$, and spins $\bm{\chi}_{1}=(-0.29,0.2,-0.3)$, $=\bm{\chi}_{2}=(-0.1,0.42,0.16)$ The $(2,2)$ mode is aligned in the early inspiral, and the time and phase shifts obtained used also to align the $(2,1)$ mode. Remarkably, the EOB and NR waveforms maintain phase coherence and a good amplitude agreement throughout the entire inspiral and merger, with the exception of the last few cycles, where the EOB model exhibits a phase lag of $\sim 1$ rad. The spin-induced precession of the system is clearly visible in the $(2,1)$ mode: nodes of the amplitude correspond to the instants at which the azimuthal angle between the inertial frame and the co-precessing frame is equal to 00. The mismatch [90] among the two waveforms is $\sim 5\times 10^{-3}$ for masses compatible with GW150914.
Figure 3: Comparison of the $(2,2)$ and $(2,1)$ modes of the TEOBResumS-Dalí model against the RIT:BBH:1632 simulation, which is an eccentric, precessing simulation with mass ratio $q=1$ and aligned spins $\chi_{1}=(0.7,0,0)$, $\chi_{2}=(0.7,0,0)$. Similarly to the case of SXS:BBH:1389, the $(2,2)$ mode is only weakly affected by the presence of precession. Eccentricity, on the other hand, induces very recognizable oscillations in the amplitude and frequency of the waveform associated to the radial motion (pericenter passages) of the binary. The mismatch [90] among the two waveforms is $\mathcal{O}(10^{-2})$ for masses compatible with GW150914.
@@ 3 GW parameter estimation
[…]
@@
We evaluate the posterior distribution $p(\bm{\theta}|d)$ using the bilby package [92], coupled with the nested sampling algorithm as implemented in dynesty [95]. We use $2048$ livepoints and $60$ as number of accepted steps (naccept), running the analysis on $48$ cores. We analyze $8$ seconds of data around the GPS time of the event, focusing on the frequencies between $20$ and $896$ Hz. We download the data from the Gravitational Wave Open Science Center (GWOSC) and use the publicly available power spectral density (PSD) and calibration envelopes as computed by the LVK collaboration [96]. We generate waveforms with TEOBResumS-Dalí using a sampling rate of $4096$ Hz, and consider the $(\ell,|m|)=(2,1),(2,2),(3,3)$ and $(4,4)$ co-precessing modes for the waveform generation. We use an initial orbit-averaged frequency for waveform generation of $15$ Hz, lower than the minimum one in the analysis. This choice is motivated by the fact that the instantaneous frequency of an eccentric waveform varies throughout each orbit. Therefore, if we were to generate waveforms starting at $20$Hz we would risk miss some signal power. As our model natively outputs waveforms in the time domain, each likelihood evaluation requires us to perform a Fast Fourier transform (FFT). With these settings, the analyses take between $9$ and $17$ days to complete, depending on the model complexity.
@@
$\log{B}^{S}_{N}$ $280.55^{+0.11}_{-0.11}$ $280.44^{+0.11}_{-0.11}$ $279.41^{+0.12}_{-0.12}$ $279.29^{+0.12}_{-0.12}$
Table 1: Summary of the parameters of GW150914 from GWTC-2.1 (first column) and under the four hypotheses considered in this work. The (natural) logarithm of the BF is reported in the last row, confirming the expectation that this event is most consistent with a quasi-circular inspiral of two BHs, and that non-circularity is disfavoured.
@@ 4 Results
When comparing the different hypotheses in terms of log Bayes factors, we do not find any scenario to be strongly favored over the others (see Tab. 1). However, we do find that adding non-circularity to the model tends to decrease its Bayesian evidence, with the QC-AS and QC-PS models being favored over their NC counterparts by $\log\mathcal{B}^{\rm QC}_{\rm NC}\sim 1$. As such, we conclude that the data is consistent with a quasi-circular BBH merger, with small effective spin and no strong evidence in favor of precession. To visually confirm our result, we show the best fitting (maximum likelihood) waveform from the non-circular, precessing-spin model in Fig. 6, overlaid on the whitened strain data from the Hanford and Livingston detectors. Compared to the simulations considered in the previous section, the (whitened) reconstructed waveform shows no clear signatures of eccentricity or precession.
@@
Figure 6: Whitened strain data from the Hanford (top) and Livingston (bottom) detectors, with the maximum likelihood waveform from the non-circular, precessing-spin model overlaid. As can be seen, the signal is visually consistent with a QC BBH merger, with no clear evidence of eccentricity or precession.
@@ Acknowledgments
The authors would like to thank S. Bernuzzi and A. Nagar for useful discussions. RG is grateful to K. Chandra and I. Gupta for constant support and encouragement throughout the development of this work, and to M. Poulsen, J. Larsen and K.B. Larsen for inspiration. The version of TEOBResumS-Dalí employed in this work is available at https://bitbucket.org/teobresums/teobresums/src/GIOTTO/ and is tagged with the arXiv number of this work. RG acknowledges support from NSF Grant PHY-2020275 (Network for Neutrinos, Nuclear Astrophysics, and Symmetries (N3AS)). JL and DC acknowledge support from the Italian Ministry of University and Research (MUR) via the PRIN 2022ZHYFA2, GRavitational wavEform models for coalescing compAct binaries with eccenTricity (GREAT). ST acknowledges support from NSF grant NSF PHY-2207780. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The authors are grateful for computational resources provided by LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459.

### v2

@@ 1 Introduction
where $|\bm{\chi}_{i}^{\perp}|$ are the projections of the spin vectors onto the plane perpendicular to $\bm{L}$. The waveform polarizations $_{+}, h×h_{\times}$ are decomposed in terms of the s=2s=-2 spin-weighted spherical harmonics $_{-2}Y_{lm}$ as
@@ 2 Modeling GWs from BBHs on generic orbits
Figure 2: Comparison of the $(2,2)$ and $(2,1)$ modes of the TEOBResumS-Dalí model against the SXS:BBH:1389 simulation, which is a spin-precessing simulation with mass ratio $q=1.63$, and spins $\bm{\chi}_{1}=(-0.29,0.2,-0.3)$, $\bm{\chi}_{2}=(-0.1,0.42,0.16)$ The $(2,2)$ mode is aligned in the early inspiral, and the time and phase shifts obtained used also to align the $(2,1)$ mode. Remarkably, the EOB and NR waveforms maintain phase coherence and a good amplitude agreement throughout the entire inspiral and merger, with the exception of the last few cycles, where the EOB model exhibits a phase lag of $\sim 1$ rad. The spin-induced precession of the system is clearly visible in the $(2,1)$ mode: nodes of the amplitude correspond to the instants at which the azimuthal angle between the inertial frame and the co-precessing frame is equal to 0. The mismatch [90] among the two waveforms is $\sim 5\times 10^{-3}$ for masses compatible with GW150914.
Figure 3: Comparison of the $(2,2)$ and $(2,1)$ modes of the TEOBResumS-Dalí model against the RIT:BBH:1632 simulation, which is an eccentric, precessing simulation with mass ratio $q=1$ and aligned spins $\chi_{1}=(0.7,0,0)$, $\chi_{2}=(0.7,0,0)$. Similarly to the case of SXS:BBH:1389, the $(2,2)$ mode is only weakly affected by the presence of precession. Eccentricity, on the other hand, induces very recognizable oscillations in the amplitude and frequency of the waveform associated to the radial motion (pericenter passages) of the binary. The mismatch [90] among the two waveforms is $\mathcal{O}(10{-2})$ for masses compatible with GW150914.
@@ 3 GW parameter estimation
In the expression above, $p(d|\bm{\theta},H)$ is the likelihood function, which contains the information about the GW signal and the noise of the detectors; $p(\bm{\theta}|H)$ is the prior distribution, which encodes our assumptions about the distribution of the parameters before the data is observed, and is usually chosen to be uniform in the parameters of interest (“agnostic” prior), or motivated by physical models; $p(d|H)=\int d\bm{\theta}p(d|\bm{\theta},H)p(\bm{\theta}|H)$ is the evidence (or “marginal likelihood”), a constant that ensures the posterior is properly normalized. In the context of model selection, $p(d|H)$ is used to compute the Bayes’ factor (BF) between two models, which quantifies the relative probability of the data given the two models and is used to determine whether one is favored by the data:
@@
We evaluate the posterior distribution $p(\bm{\theta}|d)$ using the bilby package [92], coupled with the nested sampling algorithm as implemented in dynesty [95]. We use $2048$ livepoints and $60$ as number of accepted steps (naccept), running the analysis on $48$ cores. We analyze $8$ seconds of data around the GPS time of the event, focusing on the frequencies between $20$ and $896$ Hz. We download the data from the Gravitational Wave Open Science Center (GWOSC) and use the publicly available power spectral density (PSD) and calibration envelopes as computed by the LVK collaboration [96]. We generate waveforms with TEOBResumS-Dalí using a sampling rate of $4096$ Hz, and consider the $(\ell,|m|)=(2,1),(2,2),(3,3)$ and $(4,4)$ co-precessing modes for the waveform generation. We use an initial orbit-averaged frequency for waveform generation of $15$ Hz, lower than the minimum one in the analysis. This choice is motivated by the fact that the instantaneous frequency of an eccentric waveform varies throughout each orbit. Therefore, if we were to generate waveforms starting at 2020Hz we would risk miss some signal power. As our model natively outputs waveforms in the time domain, each likelihood evaluation requires us to perform a Fast Fourier transform (FFT). With these settings, the analyses take between $9$ and $17$ days to complete, depending on the model complexity.
@@
$\log{\mathcal{B}}^{S}_{N}$ $280.55^{+0.11}_{-0.11}$ $280.44^{+0.11}_{-0.11}$ $279.41^{+0.12}_{-0.12}$ $279.29^{+0.12}_{-0.12}$
Table 1: Summary of the parameters of GW150914 from GWTC-2.1 (first column) and under the four hypotheses considered in this work. The (natural) logarithm of the BF comparing the signal to the noise hypothesis $\log\mathcal{B}^{S}_{N}$ is reported in the last row, confirming the expectation that this event is most consistent with a quasi-circular inspiral of two BHs, and that non-circularity is disfavoured.
@@ 4 Results
When comparing the different hypotheses in terms of log Bayes factors, we do not find any scenario to be strongly favored over the others (see Tab. 1). However, we do find that adding non-circularity to the model tends to decrease its Bayesian evidence, with the QC-AS and QC-PS models being favored over their NC counterparts by $\log\mathcal{B}^{\rm QC}_{\rm NC}\sim 1$. As such, we conclude that the data is consistent with a quasi-circular BBH merger, with small effective spin and no strong evidence in favor of precession. To visually confirm our result, we show the time-domain reconstruction from the non-circular, precessing-spin model in Fig. 6, overlaid on the whitened strain data from the Hanford and Livingston detectors. Compared to the simulations considered in the previous section, the (whitened) reconstructed waveforms show no clear signatures of eccentricity or precession.
@@
Figure 6: Whitened strain data from the Hanford (top) and Livingston (bottom) detectors, with the prediction from the non-circular, precessing-spin model overlaid; the thick line corresponds to the maximum likelihood waveform, the colored region is the 90% equal-tailed credible interval. As can be seen, the signal is visually consistent with a QC BBH merger, with no clear evidence of eccentricity or precession. This is reflected in the fact that the bulk of the posterior distribution for this model has low effective and precessing spin parameters, and eccentricity consistent with zero.
@@ 5 Conclusions
Nonetheless, more work is needed to further improve GW models in view of next generation space- and ground- based detectors. Beyond improvements in their accuracy, additional physical effects should be included in the waveform generation. These include higher order harmonics beyond those considered here, including $m=0$ modes which encode GW memory [98, 99, 100, 101, 102, 103, 104], improved descriptions of tidal heating and torquing [105, 106, 107], as well as phenomenological descriptions of environmental effects [108, 109, 110]. The latter, while not expected to be relevant for GW150914 [111], may be important for the analysis of future events, especially those originating from dense stellar environments [112, 113, 114, 115, 116, 117]
The increased computational cost of TEOBResumS-Dalí due to the inclusion of both eccentricity and precession is a clear limitation for the analysis of large catalogs of GW events. Various avenues can be explored to address this issue. On the waveform generation side, the use of surrogate models [118, 119, 80, 120, 121, 122, 123] can significantly speed up the generation of the waveforms. Preliminary results in this direction have been presented, for nonspinning [124, 125] and aligned-spin [126, 127] eccentric binaries. Recently, a surrogate model for aligned-spin TEOBResumS-Dalí was constructed [123], employing a further decomposition of each harmonic mode of emission into modes with monotonically increasing frequency. Alternatively, reduced order modeling techniques [118, 128, 129, 130, 122] or analytical approximations [131, 132, 133] can also be employed to speed up the generation of the waveforms. On the PE side, the use of more efficient sampling algorithms [134, 135], deep-learning techniques [11, 136] used as high-quality posterior proposals [13] or approximate or marginalized likelihoods [137, 138, 139] can significantly reduce the computational cost of the analysis; for a review, see Ref. [140].
@@ Acknowledgments
The authors would like to thank S. Bernuzzi and A. Nagar for useful discussions. RG is grateful to K. Chandra and I. Gupta for constant support and encouragement throughout the development of this work, and to M. Poulsen, J. Larsen and K.B. Larsen for inspiration. The version of TEOBResumS-Dalí employed in this work is available at https://bitbucket.org/teobresums/teobresums/src/2505.21612/. RG acknowledges support from NSF Grant PHY-2020275 (Network for Neutrinos, Nuclear Astrophysics, and Symmetries (N3AS)). JL and DC acknowledge support from the Italian Ministry of University and Research (MUR) via the PRIN 2022ZHYFA2, GRavitational wavEform models for coalescing compAct binaries with eccenTricity (GREAT). ST acknowledges support from NSF grant NSF PHY-2207780. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The authors are grateful for computational resources provided by LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459.